# Verification-Guided Shielding for Deep Reinforcement Learning

### v1

@@ Abstract
In recent years, Deep Reinforcement Learning (DRL) has emerged as an effective approach to solving real-world tasks. However, despite their successes, DRL-based policies suffer from poor reliability, which limits their deployment in safety-critical domains. As a result, various methods have been put forth to address this issue by providing formal safety guarantees. Two main approaches include shielding and verification. While shielding ensures the safe behavior of the policy by employing an external online component (i.e., a “shield”) that overruns potentially dangerous actions, this approach has a significant computational cost as the shield must be invoked at runtime to validate every decision. On the other hand, verification is an offline process that can identify policies that are unsafe, prior to their deployment, yet, without providing alternative actions when such a policy is deemed unsafe. In this work, we present verification-guided shielding — a novel approach that bridges the DRL reliability gap by integrating these two methods. Our approach combines both formal and probabilistic verification tools to partition the input domain into safe and unsafe regions. In addition, we employ clustering and symbolic representation procedures that compress the unsafe regions into a compact representation. This, in turn, allows to temporarily activate the shield solely in (potentially) unsafe regions, in an efficient manner. Our novel approach allows to significantly reduce runtime overhead while still preserving formal safety guarantees. We extensively evaluate our approach on two benchmarks from the robotic navigation domain, as well as provide an in-depth analysis of its scalability and completeness. Both authors contributed equally.
@@ 1 Introduction
Deep reinforcement learning (DRL) is gaining popularity due to its recent success in solving complex decision-making problems across various domains and settings (Mnih et al., 2013; Kober et al., 2013; Rolf et al., 2023; Karamzade et al., 2024). However, upon formal and rigorous analysis, even policies generated by state-of-the-art algorithms exhibit a significant drawback: they can not ensure the correctness of the DRL policy throughout every possible input (Katz et al., 2019b; Corsi et al., 2021). This limitation hinders the full integration of DRL agents in safety-critical scenarios, such as autonomous navigation systems (Tai et al., 2017), robotic controllers (Aractingi et al., 2023), and healthcare (Pore et al., 2021). In which even a single mistake can have dire consequences. This setback emphasizes the need for ensuring the absolute compliance of DRL policies with user-specified safety and behavioral requirements (Ray et al., 2019; Ma et al., 2024). To this end, the DRL community has recently made significant efforts to generate more reliable agents. These attempts include online training methods such as constrained optimization (Yang et al., 2022a; Zhang et al., 2020) and safe exploration (Simão et al., 2021; Kamran et al., 2022). However, despite promising results, these approaches, and others, are heuristic in nature and are unable to guarantee the absolute correctness of the DRL policy in question. This limitation is observed even when the policy is generated with state-of-the-art algorithms, and on relatively simple tasks (Corsi et al., 2024).
@@
An alternative family of approaches tackles the DRL safety problem from a different perspective, by decoupling the safety requirements from the training procedure. These techniques typically involve a formal analysis of the neural network function or the integration of various types of domain expert knowledge into the policy. Two of the most promising approaches in this context are formal verification (Liu et al., 2019) and shielding (Bloem et al., 2015). Unlike training-based methods, these techniques are indeed able to provide rigorous guarantees, however, they suffer from various limitations. Formal verification, for example, is computationally hard (Katz et al., 2017), and hence with limited applicability to various use cases (e.g., it is not clear how to verify large language models). In addition, formal verification tools typically return a binary answer, indicating whether the safety requirement holds or not, without providing any alternative solution when the latter occurs, and the policy is deemed unsafe. On the other hand, shielding techniques introduce an external component (i.e., a “shield”), that can override the original unsafe decisions, hence providing a safe action when encountering an unsafe input. However, although shielding affords safety certifications, there is still no guarantee that the proposed action is optimal (Alshiekh et al., 2018) and, crucially, the external shield must be invoked in every time-step, resulting in significant overhead. This issue is critical as, in many real-time applications, such an overhead may be unfeasible in practice.
@@
In this work, we begin bridging this gap, and present verification-guided shielding, a novel method that combines both these aforementioned techniques. Our approach consists of two main stages. First, we employ a combination of different formal methods to identify all the regions in the input space where the agent is guaranteed to behave correctly. In these regions, we can rely on the original policy, without invoking the external shield to validate the agent’s decisions (we note that this is possible only due to the rigorous guarantee provided by the formal verification process). Then, in the remaining (unsafe) input region, we activate the shield, which can potentially overrun the unsafe decision, when encountered. Our approach significantly reduces the overall overhead, while still preserving the formal guarantees regarding the policy’s safety. Implementing our approach poses several challenges, ranging from scalability to the soundness of the algorithm, which we extensively analyze in the rest of the paper. Finally, to demonstrate the effectiveness of our approach, we extensively evaluate it on two popular DRL benchmarks: (i) Particle World, where an agent is trained to navigate in a two-dimensional grid, and (ii) Mapless Navigation, a real-world task in robotics, where a robot learns to navigate in an unknown arena and reach a given target (Pore et al., 2021; Corsi et al., 2021). We use expressive shields for such tasks (Rodriguez et al., 2024).
@@
The rest of the paper is organized as follows. Sec 2 contains background on safe DRL, formal verification, and shielding. We formalize our problem in Sec 3. In Sec. 4 we present our novel method for verification-guided shielding, and empirically evaluate it in Sec. 5. Related work is covered in Sec. 6, and we conclude in Sec. 7.
@@ 2 Preliminaries
Deep reinforcement learning algorithms typically aim to optimize the expected cumulative reward, which represents the main objective of the agent (Sutton & Barto, 2018). However, in safety-critical tasks, it is common to introduce an additional function that represents the safety constraints that should be minimized as part of the optimization process. Finding a successful policy under these multiple objectives has emerged as a challenging problem (Ma et al., 2024). Moreover, it is important to note that DRL training algorithms are designated to fulfill requirements only in expectation, without any formal guarantee on the behavior of the policy during deployment.
@@ 2.1 Formal Verification
The precondition $\mathcal{P}$ usually encodes domain-specific knowledge on the input space, e.g., it can limit the inputs to represent a specific dangerous situation. The postcondition $\mathcal{Q}$ encodes the negation of the desired behavior when the agent’s current state belongs to $\mathcal{P}$. Hence, when a verification algorithm (a.k.a. “verifier”) answers UNSAT, i.e., that there does not exist a satisfying assignment, this indicates that the DNN behaves correctly on all inputs in our domain of interest. On the other hand, when the verification algorithm returns SAT, this indicates that a satisfying assignment is found, and at least a single input $x$ adheres to $\mathcal{P}(x)\wedge\mathcal{Q}(\mathcal{N}(x))$, and triggers the unwanted behavior. The DNN verification problem is computationally hard and has been proven to be NP-complete (Katz et al., 2017), hence, such techniques are afforded only in safety-critical tasks, in which the safety of the DRL in question must be rigorously guaranteed, and classic testing techniques are inadequate.
DNN Verification Example. DNN verification can be employed in many real-world problems. For instance, it has been shown that DNNs are susceptible to adversarial inputs, i.e., small input perturbations that can cause even the best DNNs to fail miserably (Szegedy et al., 2013; Huang et al., 2017a; Ma et al., 2020; Ferhat & Yildirim-Yayilgan, 2020; Gongye et al., 2020). The resilience, or robustness, of a DNN to such perturbations can directly be assessed using off-the-shelf verifiers (Tjeng et al., 2017; Zhang et al., 2018; Gopinath et al., 2018; Casadio et al., 2022) by encoding as a precondition an $\epsilon$-ball around a given input $x$ ($\mathcal{P}(x)\coloneqq B_{\epsilon}(x)$), and a postcondition as a case in which the DNN misclassifies a given input $x^{\prime}\in B_{\epsilon}(x)$. In the context of deep reinforcement learning, the verified properties are typically safety constraints, that are also encoded (by domain experts) as input-output relations. For additional details, see Appendix D.
@@ 2.2 LTL Synthesis and Shielding
DRL Shielding. Recently, it has been shown that a given LTL formula $\varphi$ represents a desired specification that can be used to automatically synthesize shields (Bloem et al., 2015; Alshiekh et al., 2018), i.e., generate external components that are coupled with the agent, and force it to behave safely according to the specification $\varphi$. More formally, given an LTL specification $\varphi$, it is possible to generate a shield $S$ w.r.t. a given system $D$ (controlled by a DRL agent, in our case). $S$ guarantees that all behaviors of the DRL-controlled
@@
system $D$ satisfy $\varphi$ as follows: when $S$ encounters an input $I$ that triggers an erroneous output (i.e., $D(I)\coloneqq O$ for which $\varphi(I,O)$ does not hold), the original action $O$ is corrected, and replaced with another action $O^{\prime}$, ensuring that $\varphi$(I, O’) does hold. This scheme, depicted in Fig. 1, guarantees that the combined system $D\cdot S$ never violates $\varphi$.
@@
Figure 2: The environments analyzed in our evaluation: Particle World (left) and Mapless Navigation (right). The graph displays the number of episodes on the x-axis and the corresponding success rate on the y-axis.
@@ 3 Motivation, Benchmarks, and Problem Formulation
Benchmarks. In our evaluation, we focused on two popular DRL benchmarks: (i) Particle World, in which an agent moves in a simple two-dimensional grid trying to reach a target position while avoiding collisions with obstacles; and (ii) Mapless Navigation, a real-world robotic navigation task, in which a robotic agent navigates in un unknown arena by relying only on local sensors. Both benchmarks are extensively studied in the context of safe DRL (Marchesini et al., 2022; Amir et al., 2023a; Corsi et al., 2024) given their straightforward safety requirements (e.g., collision avoidance). For a more detailed description of the environment and training setup, we refer to Appendix A.
@@
Motivation. All five models attained an average success rate of approximately $95\%$, and also a (seemingly) safe decision-making policy: in $100$ randomly generated trials, not a single collision was recorded. However, when analyzed through the lens of formal verification, we identified that all the selected models had input configurations in which the policies can indeed behave unsafely and collide (see the rightmost column of Tab. 1). We believe this further motivates our work — although the models were extensively trained with a s.o.t.a. algorithm on a (relatively) simple task, and although they seemed to behave safely, this was indeed not the case, as formal methods were able to uncover input configurations in which they fail miserably.
@@ 3.1 Model Selection via Verification
The well-known susceptibility of DNNs in general, and DRL agents in particular, to adversarial inputs renders it unlikely to find models that always behave safely across the whole input domain, even when trained for relatively simple tasks (Casadio et al., 2022; Amir et al., 2023a; Corsi et al., 2024). This phenomenon has also been confirmed by our evaluation reported in Tab. 1: even when formally analyzing hundreds of trained models with near-perfect empirical performance, not a single model was found to be safe throughout the entire input domain. Moreover, as discussed in Sec. 2, formal verification algorithms can detect (offline) whether a DNN is unsafe or not, however, once a DNN is deemed unsafe, it is not clear what practitioners should do. We believe, that both aforementioned limitations further motivate the need for shielding.
@@
Table 1: Results of the formal analysis for the top five models trained on the Particle World environment.
Seed No Shield Collisions (%) Shield Collisions (%) Interventions (%) Overhead
@@
Table 2: Overhead due to the standard shield; the first two columns demonstrate how the shield can prevent collisions while introducing a significant overhead, even though less than $9\%$ of the actions are overridden on average. Data is collected from the Particle World environment.
@@ 3.2 Shielding Policies in Particle Domain
Shield synthesis is the logic-based procedure of generating a shield corresponding to an LTL specification $\psi$, as explained briefly in Sec. 2. This process builds upon encoding Boolean predicates that represent the various input variables, and hence, shield synthesis is usually geared towards a finite state space. However, in most DRL use cases, including the ones covered here, there is an infinite input domain, for example, the continuous input space of Particle World. Still, it has very recently been proven that the task of synthesizing shields for such cases (formally known as LTL modulo first-order theories), is decidable (Rodriguez & Sánchez, 2023; 2024b) for various cases pertaining to the temporal logic encodings of $\psi$. Building upon these results, (Rodriguez & Sánchez, 2024a) present a novel technique that can be leveraged to synthesize shields in such scenarios (Rodriguez et al., 2024), which we use in this paper. We also note that the shield synthesis procedure can be expedited in various cases. For example, many specifications of interest are in the form $\varphi\coloneqq\LTLsquare\varphi^{\prime}$, where $\varphi^{\prime}$ is free of temporal operators. In such cases, the synthesis process can be significantly optimized and computed with alternative runtime enforcement methods (Cassandras & Lafortune, 1999; Falcone et al., 2012). Note that the shield checks both the correctness of the original action and also provides the corrected action, when necessary.
Tab. 2 summarizes the safety of five trained DRL policies, with and without shielding. Although the shield’s soundness indeed guarantees absolute safety (see the third column indicating no collisions), our results demonstrate the main limitations of current shielding approaches: the shield is activated online in every time-step throughout the whole input domain, resulting to a significant overhead during deployment. From our preliminary analysis, we found that in most cases the shield does not override the original action (i.e., no interventions; see fourth column), demonstrating the safety of the original behavior in the majority of cases. However, even in such situations, the overhead remains due to the constant shield execution. In the following section, we propose an approach that takes advantage of this and drastically reduces the number of calls to the shield while still guaranteeing safety throughout the entire input domain.

### v2

@@ Abstract
In recent years, Deep Reinforcement Learning (DRL) has emerged as an effective approach to solving real-world tasks. However, despite their successes, DRL-based policies suffer from poor reliability, which limits their deployment in safety-critical domains. Various methods have been put forth to address this issue by providing formal safety guarantees. Two main approaches include shielding and verification. While shielding ensures the safe behavior of the policy by employing an external online component (i.e., a “shield”) that overrides potentially dangerous actions, this approach has a significant computational cost as the shield must be invoked at runtime to validate every decision. On the other hand, verification is an offline process that can identify policies that are unsafe, prior to their deployment, yet, without providing alternative actions when such a policy is deemed unsafe. In this work, we present verification-guided shielding — a novel approach that bridges the DRL reliability gap by integrating these two methods. Our approach combines both formal and probabilistic verification tools to partition the input domain into safe and unsafe regions. In addition, we employ clustering and symbolic representation procedures that compress the unsafe regions into a compact representation. This, in turn, allows to temporarily activate the shield solely in (potentially) unsafe regions, in an efficient manner. Our novel approach allows to significantly reduce runtime overhead while still preserving formal safety guarantees. We extensively evaluate our approach on two benchmarks from the robotic navigation domain, as well as provide an in-depth analysis of its scalability and completeness. Both authors contributed equally.
@@ 1 Introduction
Deep reinforcement learning (DRL) is gaining popularity due to its recent success in solving complex decision-making problems across various domains and settings (Mnih et al., 2013; Kober et al., 2013; Rolf et al., 2023; Karamzade et al., 2024). However, upon formal and rigorous analysis, even policies generated by state-of-the-art algorithms exhibit a significant drawback: they can not ensure the correctness of the DRL policy for every possible input (Katz et al., 2019b; Corsi et al., 2021). This limitation hinders the full integration of DRL agents in safety-critical scenarios, such as autonomous navigation systems (Tai et al., 2017), robotic controllers (Aractingi et al., 2023), healthcare (Pore et al., 2021), and decision support in regulated industries (Singh et al., 2022) in which even a single mistake can have dire consequences. This setback emphasizes the need for ensuring the absolute compliance of DRL policies with user-specified safety and behavioral requirements (Ray et al., 2019; Ma et al., 2024). To this end, the DRL community has recently made significant efforts to generate more reliable agents. These attempts include online training methods such as constrained optimization (Yang et al., 2022a; Zhang et al., 2020) and safe exploration (Simão et al., 2021; Kamran et al., 2022). However, despite promising results, these approaches are heuristic in nature and are unable to guarantee the absolute correctness of the DRL policy in question. This limitation is observed even when the policy is generated with state-of-the-art algorithms, and for performing relatively simple tasks (Corsi et al., 2024).
@@
An alternative family of approaches tackles the DRL safety problem from a different perspective, by decoupling the safety requirements from the training procedure. These techniques typically involve a formal analysis of the neural network function or the integration of various types of domain expert knowledge into the policy. Two of the most promising approaches in this context are formal verification (Liu et al., 2019) and shielding (Bloem et al., 2015). Unlike training-based methods, these techniques are indeed able to provide rigorous guarantees, but they suffer from various limitations. Formal verification, for example, is computationally hard (Katz et al., 2017), and its applicability to various use cases is thus limited (e.g., it is not clear how to verify large language models). In addition, formal verification tools typically return a binary answer, indicating whether the safety requirement holds or not, without providing any alternative solution when the latter occurs, and the policy is deemed unsafe. On the other hand, shielding techniques introduce an external component (i.e., a “shield”), that can override the original unsafe decisions, hence providing a safe action when encountering an unsafe input. However, although shielding affords safety certifications, there is still no guarantee that the proposed action is optimal (Alshiekh et al., 2018) and, crucially, the external shield must be invoked in every time step, resulting in significant overhead. This issue is critical, because in many real-time applications such an overhead may be infeasible in practice.
@@
In this work, we begin bridging this gap, and present verification-guided shielding, a novel method that combines both these aforementioned techniques. Our approach consists of two main stages. First, we employ a combination of different formal methods to identify all the regions in the input space where the agent is guaranteed to behave correctly. In these regions, we can rely on the original policy, without invoking the external shield to validate the agent’s decisions (we note that this is possible only due to the rigorous guarantee provided by the formal verification process). Then, in the remaining (unsafe) input region, we activate the shield, which can potentially override the unsafe decision, when encountered. Our approach significantly reduces the overall overhead of “traditional shielding”, while still preserving the formal guarantees regarding the policy’s safety. Implementing our approach poses several challenges, ranging from scalability to the soundness of the algorithm, which we thoroughly analyze in the following sections. Finally, to demonstrate the effectiveness of our approach, we extensively evaluate it on two popular DRL benchmarks: (i) Particle World, where an agent is trained to navigate in a two-dimensional grid, and (ii) Mapless Navigation, a real-world task in robotics, where a robot learns to navigate in an unknown arena and reach a given target (Pore et al., 2021; Corsi et al., 2021). We use expressive shields for such tasks (Rodriguez et al., 2024).
@@
The rest of the paper is organized as follows. Sec. 2 contains background on safe DRL, formal verification, and shielding. We formalize our problem in Sec. 3. In Sec. 4, we present our novel method for verification-guided shielding, and empirically evaluate it in Sec. 5. Related work is covered in Sec. 6, and we conclude in Sec. 7.
@@ 2 Preliminaries
Deep reinforcement learning algorithms typically aim to optimize the expected cumulative reward, which represents the main objective of the agent (Sutton & Barto, 2018). However, in safety-critical tasks, it is common to introduce an additional function that represents the safety constraints that should be met as part of the optimization process. Finding a successful policy under these multiple objectives has emerged as a challenging problem (Ma et al., 2024). Moreover, it is important to note that DRL training algorithms are designed to fulfill requirements only in expectation, without any formal guarantee on the behavior of the policy during deployment.
@@ 2.1 Formal Verification
The precondition $\mathcal{P}$ usually encodes domain-specific knowledge on the input space, e.g., it can limit the inputs to represent a specific dangerous situation. The postcondition $\mathcal{Q}$ encodes the negation of the desired behavior when the agent’s current state belongs to $\mathcal{P}$. Hence, when a verification algorithm (the “verifier”) answers UNSAT, i.e., that there does not exist a satisfying assignment, this indicates that the DNN behaves correctly on all inputs in our domain of interest. On the other hand, when the verification algorithm returns SAT, this indicates that a satisfying assignment is found, and at least a single input $x$ adheres to $\mathcal{P}(x)\wedge\mathcal{Q}(\mathcal{N}(x))$, and triggers the unwanted behavior. The DNN verification problem is computationally hard and has been proven to be NP-complete (Katz et al., 2017), hence, such techniques are usually applied only in safety-critical tasks, in which the safety of the DRL in question must be rigorously guaranteed, and classic testing techniques are inadequate.
DNN Verification Example. DNN verification can be employed in many real-world problems. For instance, it has been shown that DNNs are susceptible to adversarial inputs, i.e., small input perturbations that can cause even the best DNNs to fail miserably (Szegedy et al., 2013; Huang et al., 2017a; Ma et al., 2020; Ferhat & Yildirim-Yayilgan, 2020; Gongye et al., 2020). The resilience, or robustness, of a DNN to such perturbations can directly be assessed using off-the-shelf verifiers (Tjeng et al., 2017; Zhang et al., 2018; Gopinath et al., 2018; Casadio et al., 2022) by encoding as a precondition an $\epsilon$-ball around a given input $x$ ($\mathcal{P}(x)\coloneqq x\in B_{\epsilon}(x)$), and as a postcondition a case in which the DNN misclassifies a given input $x^{\prime}\in B_{\epsilon}(x)$. In the context of deep reinforcement learning, the verified properties are typically safety constraints, that are also encoded (by domain experts) as input-output relations. For additional details, see Appendix D.
@@ 2.2 LTL Synthesis and Shielding
DRL Shielding. Recently, it has been shown that a given LTL formula $\varphi$ represents a desired specification that can be used to automatically synthesize shields (Bloem et al., 2015; Alshiekh et al., 2018), i.e., generate external components that are coupled with the agent, and force it to behave safely according to the specification $\varphi$. More formally, given an LTL specification $\varphi$,
@@
it is possible to generate a shield $S$ with respect to a given system $D$ (controlled by a DRL agent, in our case). $S$ guarantees that all behaviors of the DRL-controlled system $D$ satisfy $\varphi$ as follows: when $S$ encounters an input $I$ that triggers an erroneous output (i.e., $D(I)\coloneqq O$ for which $\varphi(I,O)$ does not hold), the original action $O$ is corrected, and replaced with another action $O^{\prime}$, ensuring that $\varphi$(I, O’) does hold. This scheme, depicted in Fig. 1, guarantees that the combined system $D\circ S$ never violates $\varphi$.
@@
Figure 2: The environments analyzed in our evaluation: Particle World (left) and Mapless Navigation (right). The plot displays the number of episodes on the x-axis and the corresponding success rate on the y-axis.
@@ 3 Motivation, Benchmarks, and Problem Formulation
Benchmarks. In our evaluation, we focused on two popular DRL benchmarks: (i) Particle World, in which an agent moves in a simple two-dimensional grid trying to reach a target position while avoiding collisions with obstacles; and (ii) Mapless Navigation, a real-world robotic navigation task, in which a robotic agent navigates in an unknown arena by relying only on local sensors. Both benchmarks are extensively studied in the context of safe DRL (Marchesini et al., 2022; Amir et al., 2023a; Corsi et al., 2024) given their straightforward safety requirements (e.g., collision avoidance). For a more detailed description of the environment and training setup, see Appendix A.
@@
Motivation. All five models attained an average success rate of approximately $95\%$, and also a (seemingly) safe decision-making policy: in $100$ randomly generated trials, not a single collision was recorded. However, when analyzed through the lens of formal verification, we identified that all the selected models had input configurations in which the policies can indeed behave unsafely and collide with a wall (see the rightmost column of Tab. 1). We believe this further motivates our work — although the models were extensively trained with a state-of-the-art algorithm to solve a (relatively) simple task, and although they seemed to behave safely in empirical evaluation, this was indeed not the case, as formal methods were able to uncover input configurations in which they fail miserably.
@@ 3.1 Model Selection via Verification
The well-known susceptibility of DNNs in general, and DRL agents in particular, to adversarial inputs renders it unlikely to find models that always behave safely across the whole input domain, even when trained for relatively simple tasks (Casadio et al., 2022; Amir et al., 2023a; Corsi et al., 2024). This phenomenon has also been confirmed by our evaluation reported in Tab. 1: even when formally analyzing hundreds of trained models with near-perfect empirical performance, not a single model was found to be safe throughout the entire input domain. Moreover, as discussed in Sec. 2, formal verification algorithms can detect (offline) whether a DNN is unsafe or not, however, once a DNN is deemed unsafe, it is not clear what practitioners should do. We believe that both aforementioned limitations further motivate the need for shielding.
@@
Table 1: Results of the formal analysis for the top five models trained on the Particle World environment. A SAT verification output indicates the existence of unsafe behaviors.
Seed w/o Shield Collisions (%) w/ Shield Collisions (%) Interventions (%) Overhead
@@
Table 2: Overhead due to standard shielding; the first two columns demonstrate how the shield can prevent collisions while introducing a significant overhead, even though fewer than $9\%$ of the actions are overridden on average across seeds. Data is collected from the Particle World environment.
@@ 3.2 Shielding Policies in Particle Domain
Shield synthesis is the logic-based procedure of generating a shield corresponding to an LTL specification $\varphi$, as explained briefly in Sec. 2. This process builds upon encoding Boolean predicates that represent the various input variables, and hence, shield synthesis is usually geared towards a finite state space. However, in most DRL use cases, including the ones covered here, there is an infinite input domain, for example, the continuous input space of Particle World. Still, it has very recently been proven that the task of synthesizing shields for such cases (formally known as LTL modulo first-order theories), is decidable (Rodriguez & Sánchez, 2023; 2024b) for various cases pertaining to the temporal logic encodings of $\varphi$. Building upon these results, Rodriguez & Sánchez (2024a) presents a novel technique that can be leveraged to synthesize shields in such scenarios (Rodriguez et al., 2024), which we use in this paper. We also note that the shield synthesis procedure can be expedited in various cases. For example, many specifications of interest are in the form $\varphi\coloneqq\LTLsquare\varphi^{\prime}$, where $\varphi^{\prime}$ is free of temporal operators. In such cases, the synthesis process can be significantly optimized and computed with alternative runtime enforcement methods (Cassandras & Lafortune, 1999; Falcone et al., 2012). Note that the shield both checks the correctness of the original action and provides the corrected action, when necessary.
Tab. 2 summarizes the safety of five trained DRL policies, with and without shielding. Although the shield’s soundness indeed guarantees absolute safety (see the third column indicating no collisions), our results demonstrate the main limitations of current shielding approaches: the shield is activated online in every time-step throughout the whole input domain, resulting in a significant overhead during deployment. From our preliminary analysis, we found that in most cases the shield does not override the original action (i.e., no interventions; see fourth column), demonstrating the safety of the original behavior in the majority of cases. However, even in such situations, the overhead remains due to the constant shield execution. In the following section, we propose an approach that takes advantage of this and drastically reduces the number of calls to the shield while still guaranteeing safety throughout the entire input domain.